OpenStudy (deercult):

Expand the logarithmic expression. log_b (sqrt)57/74

OpenStudy (deercult):

\[\log _{b}\sqrt{\frac{ 57 }{ 74 }}\] Here, this looks a little nicer.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

write with exponents first \[\log_b\left(\frac{57}{74}\right)^{\frac{1}{2}}\]

OpenStudy (deercult):

Hullo ;w; Ok!

OpenStudy (misty1212):

then the one half comes out front since \[\log(x^n)=n\log(x)\]

OpenStudy (misty1212):

one you do that, use the fact that \[\log(\frac{A}{B})=\log(A)-\log(B)\]

OpenStudy (misty1212):

make sure to use parentheses, because the whole thing gets multiplied by one half

OpenStudy (deercult):

So, is it: \[\frac{ 1 }{ 2 }\log _{b}57 - \frac{ 1 }{ 2 }\log_{b}74\] or \[\frac{ 1 }{ 2 }\log_{b}(57-74)\] ? Just to be sure

OpenStudy (deercult):

Sorry, Im super tired ;<;

OpenStudy (misty1212):

first one

OpenStudy (deercult):

Thank you!

OpenStudy (misty1212):

yw btw the second one makes no sense that would mean \[57\div 74=57-74\]

OpenStudy (misty1212):

\[\color\magenta \heartsuit\]

OpenStudy (deercult):

Ah, ok. Thank you for explaining!